OpenStudy (anonymous):

What is the sum of the geometric series in which a1 = 2, r = 3, and an = 486? Sn = −1,254 Sn = −728 Sn = 81 Sn = 728

OpenStudy (anonymous):

only by guessing a1=2,r=3,an=486 all are positive so sn can't be negative first two options are deleted. a1=2,an=486 so sum should be >486 so sn=728 but it is only a guess. now we will proceed systematically. \[an=a1 r^{n-1},486=2*3^{n-1},3^{n-1}=\frac{ 486 }{ 2 }=243=3^5\] n-1=5 n=5+1=6 \[Sn=a1\frac{ r^{n} -1}{ r-1 }=2\frac{ 3^6-1 }{ 3-1 }=2\frac{ 729-1 }{ 2 }=728\]

OpenStudy (anonymous):

That's the answer that I got! Thanks! I have a few more, do you think you will be able to help me?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

yes ,i will try.